OpenStudy (nick1232):

39+n=-8+2(1+5n) Last question , please help

OpenStudy (nick1232):

>.<

OpenStudy (farmdawgnation):

It looks like you've gotten a lot of help on thse so far tonight. Perhaps you'd care to enlighten us by sharing the work you've done so far?

OpenStudy (nick1232):

what do you mean by that doe ?

OpenStudy (anonymous):

39+n=-8+2(1+5n) 39+8= -n+2+10n 47= 9n +2 47 - 2 = 9n 45=9n n=45/9 or n = 5

OpenStudy (nick1232):

Thank you !

OpenStudy (nick1232):

This makes more sense now

OpenStudy (farmdawgnation):

@rushal If you'd looked in this user's history, you would have seen that they've had more than a handful of these problems solved exactly as you have solved it, and appear to still not entirely understand the problem. In the future, please work with them as I was trying to do before just spitting out a full solution. Thanks.

OpenStudy (nick1232):

That would be nice. I do the best i can to try to figure out how they did it.

OpenStudy (nick1232):

I stay after school with my teacher all the time too so im understanding it more and more

OpenStudy (farmdawgnation):

I'm going to copy the answer you were given and add some explanation.

OpenStudy (farmdawgnation):

(I'm not going to be double checking the math so if the answer was wrong then that's a different issue. heh.)

OpenStudy (farmdawgnation):

39+n=-8+2(1+5n) Starting from this point, our first goal is to try and do two things. Get rid of those parenthesis, and then get all of the n's on one side and all of the regular numbers on the other side. So, we use distribution to get rid of the 2. We multiply 2*1 and 2*5n to get rid of the parenthesis. We then add 8 to both sides and subtract n from both sides. That yields: 39+8= -n+2+10n Now all the n's are on one side and we can start working some magic. So, we combine 39+8 into 47 and 10n and -n into 9n. that gives us: 47= 9n +2 Now, that 2 doesn't have an n. That means it's on the wrong side. Subtract it from both sides. 47 - 2 = 9n 45=9n Now all that's left is to try and get n by itself. We can do this by dividing both sides by 9. n=45/9 or n = 5 Does that help?

OpenStudy (anonymous):

hey @farmdawgnation dawg(lol) i just created my account like 1 hour ago and i am so sorry for this, i just wanted to help him. I think you would understand :)